OpenStudy (anonymous):

3. Solve the equation, show all your work: 5x + 1 = 2x + 7

OpenStudy (anonymous):

4. Solve the equation, show all your work: 7x + 3 = 5 – 2x

OpenStudy (disco619):

Which is the question, 2 or 4? Or both?

OpenStudy (anonymous):

both

OpenStudy (disco619):

Ok, i'll do only 1 for you. And leave the second one for you to do...

OpenStudy (anonymous):

okay plz!!

OpenStudy (disco619):

\[5x + 1= 2x + 7\] \[5x - 2x + 1 = 7\] \[3x + 1 = 7\] \[3x = 7 - 1\] \[3x = 6\] \[x = \frac{ 6 }{ 3 }\] \[x = 2 [ANS]\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (disco619):

Can you do the Second one yourself?

OpenStudy (anonymous):

i'll try

OpenStudy (disco619):

let me know if you can't

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

i got a 5. wait yeah i have no idea what im doing.

OpenStudy (anonymous):

disco?

OpenStudy (disco619):

LoL, should i so the second part?

OpenStudy (disco619):

But this time i'll explain step by step....

OpenStudy (anonymous):

plzz, please im sorry im so bad in math

OpenStudy (disco619):

Don't worry. Im worse xD

OpenStudy (anonymous):

5x+1 = 2x +7 5x-2x = 7 - 1 3x = 6 x = 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

looking i wouldnt say hearing

OpenStudy (disco619):

Dude! Don't interrupt what Im doing. It's bad manners...

OpenStudy (disco619):

@countrygirlnakali i'll show you *how* to do the equation...

OpenStudy (disco619):

I was talking to @sprmgovind

OpenStudy (anonymous):

yeah just saw

OpenStudy (anonymous):

hey disco what happen @Disco619

OpenStudy (anonymous):

its alright sprmgovind just that he was going to help me.

OpenStudy (anonymous):

@countrygirlnakali did u got the answer

OpenStudy (anonymous):

i did, but i need more explaination to get it thanks though. :)

OpenStudy (anonymous):

sorry disco keep typing. i want to understand it sorry for more interption

OpenStudy (anonymous):

Hey you still there?